AnimeWeebHAHA:

Can someone tell me what is (56 × 8) - (9 + 34) ? Help

AnimeWeeebLMAO:

is it rounding or just regular?

Astro:

First you want to start off with..56*8=? What does that equal?

AnimeWeebHAHA:

Rounding,

AnimeWeebHAHA:

Oh ok

sxdsouls:

just simplify bruther

sxdsouls:

or evaluate

AnimeWeeebLMAO:

It's easy so since you said your rounding, multiply 60 x 10 to start off

sxdsouls:

\[(56 × 8) ~~~~ -~~~~ (9 + 34) \] vv^ multiply \[448−(9+34)\] < take that from up there ^ add 9+34 \[448−43\] so wut do u get

sxdsouls:

\(\color{red}{answer~me~now~peasant~}\)

sxdsouls:

are you not oging to answer me

v6mpiresk8ter:

LOL

sxdsouls:

are we going to have to kill eachtother cutely

AngryPomeranian:

lol

sxdsouls:

bestie i just answered it stop trynna copy me

AnimeWeeebLMAO:

He or she said it was rounding.

sxdsouls:

wut

AnimeWeeebLMAO:

\[(60 x 10) -(10 + 30) = answer \it\]

AnimeWeebHAHA:

620?

sxdsouls:

no

sxdsouls:

448−43 `-` SUBTRACTION \[~~~448\]-\[~~~~~43\] \ 8 - 3 = 5 4 - 4 = 0 4 - 0 = 4 .... 405.. @laylalyssa